OpenStudy (anonymous):

Solve the equation: the square root of 6x+8=7x-6

OpenStudy (anonymous):

Hi, is the square root applicable for 6x+8 only?

OpenStudy (anonymous):

\[\sqrt{6x+8=}\]

OpenStudy (anonymous):

\[\sqrt{7x-6}\]

OpenStudy (anonymous):

Is that elmo?

OpenStudy (anonymous):

\[4\sqrt{6+5\sqrt{6}}\]

OpenStudy (anonymous):

\[4\sqrt{6 + 5 \sqrt{6}}\]